OpenStudy (anonymous):

how do I simplify this \[\sqrt(x-2) \sqrt(1+\frac{9}{4(x-2)})\]

OpenStudy (anonymous):

the second part is all under the radical sign

OpenStudy (anonymous):

Is it? \[\sqrt{(x-2)} \sqrt{(1+\frac{9}{4(x-2)})}\]

OpenStudy (anonymous):

yes, how do you get your radical signs to extend that far?

OpenStudy (anonymous):

after sqrt use the curli brackets: {}

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Firstly solve the expression in second radical..

OpenStudy (anonymous):

can't you multiply inside the radical?

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\]

OpenStudy (anonymous):

\[= \sqrt{(x-2)(1 + \frac{9}{4(x-2)}})\]

OpenStudy (anonymous):

\[= \sqrt{(x-2)+ \frac{9}{4}}\] hey @satellite73 i have a question for you regarding simpson's rule a little later

OpenStudy (anonymous):

or now if i may ask

OpenStudy (anonymous):

a little later will have to be tomorrow, because i have got to go but post, there are lots here who can answer

OpenStudy (anonymous):

I posted it earlier and got some responses but it kinda went over my head

OpenStudy (anonymous):

how large should n be to guarantee that the Simpson's Rule approximation to ∫10ex2dx is accurate to within 0.00001?